OpenStudy (anonymous):

Solve the systems using substitution 5x - 4y = 14 x - 2y = -2

OpenStudy (anonymous):

5x - 4y = 14 =>5(2y-2) - 4y = 14 =>10y-10-4y=14 =>6y=24 =>y=4 =>x=6